# HippoMM: Hippocampal-inspired Multimodal Memory for Long Audiovisual Event Understanding

### v1

@@ 4.2 Memory Formation
4.2.1 Temporal Pattern Separation
Continuous experience is often perceived and remembered as discrete events [baldassano2017discovering]. To computationally mimic the hippocampus’s ability to segment experience (a function often linked to the dentate gyrus [yassa2011pattern]), HippoMM first employs content-adaptive temporal segmentation as a functional approximation. This stage analyzes the incoming raw audiovisual stream to identify boundaries corresponding to significant perceptual shifts, partitioning the stream into potential event segments. This contrasts with fixed-duration chunking, aiming instead for more semantically coherent and efficiently processed units. A segment boundary is indicated at time $t$ by $\mathcal{S}_{t}=1$ if a significant change is detected in either the visual or auditory stream, using computationally efficient metrics applied directly to the input:
$\mathcal{S}_{t}=\bigvee_{m\in\{v,a\}}\text{1}\left[d_{m}(\text{Input}_{m}(t),\text{Input}_{m}(t-1))>\tau_{m}\right]$ (4)
where $d_{m}$ is a modality-specific distance function comparing relevant inputs for modality $m$ at times $t$ and $t-1$, $\tau_{m}$ is a predefined threshold, $\text{1}[\cdot]$ is the indicator function, and $\bigvee$ denotes logical OR. Specifically, for the visual modality ($m=v$), we define the distance $d_{v}$ based on the Structural Similarity Index Measure (SSIM) between consecutive raw visual frames $F_{t},F_{t-1}$:
$d_{v}(F_{t},F_{t-1})=1-\text{SSIM}(F_{t},F_{t-1})$ (5)
A value close to 1 signifies substantial visual change. For the auditory modality ($m=a$), the distance $d_{a}$ detects silence or low-energy periods directly from the raw audio signal $a_{t}$, often corresponding to semantic breaks:
$d_{a}(a_{t})=-20\log_{10}\left(\sqrt{\frac{1}{N}\sum_{i=1}^{N}a_{i}^{2}}\right)$ (6)
where $a_{i}$ are audio samples within a window and $N$ is the window size. High $d_{a}$ values indicate silence. While algorithmically distinct from neural sparse coding, this heuristic boundary detection serves the functional goal of enhancing temporal discriminability. Segment lengths are constrained between $t_{min}$ and $t_{max}$ via merging or splitting based on these detected boundaries. This segmentation process defines the temporal windows $[t_{start,i},t_{end,i}]$ for subsequent detailed encoding and memory storage.
4.2.2 Perceptual Encoding
Building upon the temporal segmentation which delineates the continuous stream into segments $i$ defined by boundaries $[t_{start,i},t_{end,i}]$, the Perceptual Encoding stage processes the raw multimodal input $x_{t}$ within each segment. Its goal is transforming the input into rich, multimodal representations suitable for memory storage and recall, mirroring the entorhinal cortex’s role in relaying sensory information to the hippocampus using powerful pre-trained models. For any given time step $t$ within a segment $i$, the system applies a tripartite encoding strategy. The visual pathway ($v$) processes $x_{t}^{v}$ using a vision-language model for descriptions $\mathcal{T}_{v}(x_{t}^{v})$. The auditory pathway ($a$) processes $x_{t}^{a}$ via speech recognition for transcriptions $\mathcal{T}_{a}(x_{t}^{a})$. The cross-modal pathway ($c$) uses a joint embedding model on $x_{t}^{v}$ and $x_{t}^{a}$ for a unified feature vector $\mathcal{E}_{c}(x_{t}^{v},x_{t}^{a})$, with optional normalization $\phi_{c}$. The collective instantaneous output is the multimodal trace $\mathcal{M}_{t}=\{\mathcal{T}_{v}(x_{t}^{v}),\mathcal{T}_{a}(x_{t}^{a}),\mathcal{E}_{c}(x_{t}^{v},x_{t}^{a})\}$.
These traces $\mathcal{M}_{t}$ are aggregated over the duration of segment $i$ (for $t$ from $t_{start,i}$ to $t_{end,i}$) to form segment-level representations: the sequence of cross-modal embeddings $\mathbf{E}_{i}=\{\mathcal{E}_{c}(x_{k}^{v},x_{k}^{a})\mid k\in[t_{start,i},t_{end,i}]\}$ and the textual descriptions $\mathbf{T}_{i}=\{\mathcal{T}_{v}(x_{k}^{v}),\mathcal{T}_{a}(x_{k}^{a})\mid k\in[t_{start,i},t_{end,i}]\}$. This aggregated content is then stored within a ShortTermMemory object $m_{i}$, which also holds pointers $\mathbf{C}_{i}$ to the raw audiovisual snippets and segment timestamps $t_{i},s_{i}$:
$m_{i}=\{\mathbf{E}_{i},\mathbf{T}_{i},\mathbf{C}_{i},t_{i},s_{i}\}$ (7)
This object $m_{i}$, representing the processed segment, serves as the short-term memory for the subsequent stage.
4.2.3 Memory Consolidation
Inspired by biological memory consolidation processes that stabilize and optimize memory traces [rolls2013quantitative], this stage aims to reduce redundancy in the sequence of ShortTermMemory objects $\{m_{i}\}$ generated by the Perceptual Encoding stage. Storing highly similar consecutive segments $m_{i},m_{i+1}$ can be inefficient. We implement a filtering mechanism based on semantic similarity in the shared cross-modal embedding space. For each segment $m_{i}$, we derive a representative cross-modal embedding $\mathbf{v}_{i}$ by averaging the sequence of embeddings $\mathbf{E}_{i}$ stored within $m_{i}$. We then maintain a set $K$ of indices corresponding to the retained, non-redundant segments. A segment $i$ is added to $K$ only if its representative feature vector $\mathbf{v}_{i}$ is sufficiently dissimilar from the representative vectors $\mathbf{v}_{j}$ of already retained segments $j$ ($j\in K,j<i$), assessed via a cosine similarity threshold $\gamma$:
$K=\{i\mid\forall j\in K,j<i\implies\cos(\mathbf{v}_{i},\mathbf{v}_{j})<\gamma\}$ (8)
Segments $i\notin K$ (i.e., those where $\cos(\mathbf{v}_{i},\mathbf{v}_{j})\geq\gamma$ for some $j\in K$) are considered redundant and discarded or merged. This sparsification reduces storage and computational load during retrieval while preserving segments that introduce significant new information, functionally analogous to interference reduction mechanisms in biological memory. The output of this stage is the sequence of consolidated ShortTermMemory objects $\{m_{k}\mid k\in K\}$.
4.2.4 Semantic Replay
Following consolidation, HippoMM implements a crucial stage inspired by neuroscientific concepts of memory replay and abstraction [eichenbaum2017prefrontal], serving as the core mechanism for transforming detailed, consolidated short-term memories into efficient, semantic long-term representations. This Semantic Replay process operates on the sequence of consolidated ShortTermMemory objects $\{m_{k}\mid k\in K\}$. Functionally, it “replays” the multimodal essence of each detailed segment $m_{k}$ through an LLM, denoted $\phi_{\text{LLM}}$, to extract and synthesize its core semantic meaning or “gist”, analogous to how hippocampal replay during sleep consolidates daytime experiences into integrated semantic knowledge in biological systems.
To generate this abstract representation, the LLM integrates key multimodal information derived from the corresponding detailed segment $m_{k}$. Specifically, the LLM processes structured multimodal context derived from $m_{k}$, including key visual information $\mathbf{I}_{\text{visual}}(m_{k})$ (extracted from visual descriptions in $\mathbf{T}_{k}$ and potentially keyframes referenced by $\mathbf{C}_{k}$) and relevant auditory information $\mathbf{A}_{\text{audio}}(m_{k})$ (extracted from transcriptions in $\mathbf{T}_{k}$), formatted to facilitate semantic summarization:
$\mathbf{S}_{\theta_{k}}=\phi_{\text{LLM}}(\text{FormattedContext}(m_{k}))$ (9)
The resulting textual summary $\mathbf{S}_{\theta_{k}}$ encapsulates the high-level semantic content distilled from the detailed segment $m_{k}$.
This semantic summary forms the centerpiece of a ThetaEvent object $\theta_{k}$, which constitutes the abstracted long-term memory representation corresponding to the detailed short-term memory $m_{k}$:
$\theta_{k}=\{\mathbf{v}_{k},\mathbf{T}_{\theta_{k}},\mathbf{I}_{\theta_{k}},\mathbf{A}_{\theta_{k}},\mathbf{S}_{\theta_{k}}\}$ (10)
Each ThetaEvent object $\theta_{k}$, representing an abstracted long-term memory representation, bundles the representative cross-modal embedding $\mathbf{v}_{k}$ (identical to the one computed during consolidation), the segment’s defining temporal information $\mathbf{T}_{\theta_{k}}$ (e.g., start/end timestamps derived from $m_{k}$’s $t_{k},s_{k}$), references to key visual ($\mathbf{I}_{\theta_{k}}$) and auditory ($\mathbf{A}_{\theta_{k}}$) context used during summarization, and the core semantic summary $\mathbf{S}_{\theta_{k}}$. This crucial temporal information $\mathbf{T}_{\theta_{k}}$ provides grounding and serves as the linkage mechanism back to the detailed consolidated ShortTermMemory object $m_{k}$, enabling transitions between abstract and detailed recall. The collection of these ThetaEvent objects $\{\theta_{k}\}$ constitutes HippoMM’s abstracted long-term memory store, facilitating efficient gist-based semantic search and reasoning while complementing the detailed episodic information retained in the consolidated short-term store $\{m_{k}\}$. This explicit transformation from detailed traces ($m_{k}$) to abstract summaries ($\theta_{k}$) operationalizes a key aspect of biological short-term to long-term memory systems, enabling more scalable and flexible memory querying. The ThetaEvent name draws inspiration from hippocampal theta rhythms’ role in temporal organization and memory sequencing [buzsaki2002theta], without explicit oscillatory modeling.
@@ 4.2.4 Semantic Replay
Table 1: Performance comparison of HippoMM against existing methods and ablation variants. PT: Processing Time; ART: Average Response Time; A+V: Cross-modal (Audio+Visual) accuracy; A: Audio-only accuracy; V: Visual-only accuracy; S: Semantic understanding accuracy. Best results are in bold, second best are underlined. HippoMM significantly outperforms prior methods in all modality-specific tasks while maintaining efficient processing and response times. Ablation studies demonstrate the importance of each component: Detailed Recall (DR), Fast Retrieval (FR), and Adaptive Reasoning (AR).
@@
NotebookLM 28.40% 23.20% 28.00% 26.80% 26.60%
Video RAG 9.46h 112.5s 63.6% 67.2% 41.2% 84.8% 64.2%
@@
4.3 Memory Retrieval
4.3.1 Hierarchical Retrieval Architecture with Query Analysis
Memory retrieval commences with an analysis of the user query $q$ to discern its nature and inform subsequent retrieval steps. A query classifier, $\mathcal{Q}_{\text{type}}(q)$, determines the query’s primary modality focus (e.g., Visual $V$, Auditory $A$, Cross-modal $V\times A$) or identifies it as targeting high-level semantic/gist information. To balance retrieval efficiency and depth, HippoMM employs a hierarchical strategy that prioritizes the faster pathway. It always begins by attempting Fast Retrieval ($\Phi_{\text{fast}}$), operating on the set of abstract semantic summaries =$\mathcal{E}=\{\mathbf{S}_{\theta_{k}}\}$ derived from the ThetaEvent objects (Figure 2b, upper path). This pathway is optimized for rapidly answering semantic or gist-level queries and directly generates a candidate answer along with a confidence score. Only if the confidence score of the result from $\Phi_{\text{fast}}$ falls below a predefined threshold $\tau$, does the system escalate to the more resource-intensive Detailed Recall ($\Psi_{\text{detailed}}$). This second stage operates on the richer, consolidated ShortTermMemory objects =$\mathcal{M}=\{m_{k}\mid k\in K\}$ (Figure 2b, lower path). Crucially, should this escalation occur, the specific mechanisms employed within Detailed Recall are explicitly guided by the query type determined during the initial analysis ($\mathcal{Q}_{\text{type}}(q)$), ensuring an appropriate search strategy is used. The overall retrieval process $\mathcal{R}(q)$ follows this confidence-gated hierarchy:
$\mathcal{R}(q)=\begin{cases}\Phi_{\text{fast}}(\mathcal{E},q)&\text{if }\text{conf}(\Phi_{\text{fast}})>\tau\\ \Psi_{\text{detailed}}(q,\mathcal{M},\mathcal{Q}_{\text{type}}(q))&\text{otherwise}\end{cases}$ (11)
The hyperparameter $\tau$ solely mediates the switch between these answer-generating pathways, while the pre-determined query type directs the execution approach within $\Psi_{\text{detailed}}$ when invoked.
4.3.2 Detailed Recall Pathways
The Detailed Recall process ($\Psi_{\text{detailed}}$), invoked when Fast Retrieval yields insufficient confidence (Sec. 4.3.1), utilizes the query’s modality focus (Visual $V$, Auditory $A$, or Cross-modal $V\times A$) — identified by the classifier $\mathcal{Q}_{\text{type}}(q)$ during the initial query analysis phase — to select the appropriate search strategy within the consolidated ShortTermMemory objects $\{m_{k}\}$. The pathway chosen depends on this predetermined query type:
For single-modality queries (targeting primarily visual or auditory information), search within the relevant content of $\{m_{k}\}$ proceeds in two potential paths. First, if the query targets perceptual similarity and can be represented as an embedding $q_{embed}$, embedding-based retrieval (Feature Search) attempts a direct match using feature similarity against the representative embeddings $\{\mathbf{v}_{k}\}$ or potentially the finer-grained embedding sequences $\{\mathbf{E}_{k}\}$ stored within $\{m_{k}\}$. If this feature-based search fails, is inapplicable, or yields low confidence, text-based retrieval (Semantic Search) is performed using the aggregated textual descriptions $\mathbf{T}_{k}$ (containing captions, transcriptions) within each $m_{k}$.
For cross-modal queries (requiring association between modalities), retrieval leverages temporal co-occurrence as a proxy for episodic binding. The process unfolds in three steps: First, the system uses the pre-computed query embedding $q_{embed}$ representing the query cue. It identifies the top-$k$ segments $m_{k}$ whose representative cross-modal embedding $\mathbf{v}_{k}$ is most similar to $q_{embed}$, finding segments related to the query cue modality:
$\mathbf{S}_{\text{query}}$ $=\text{TopK}(\text{sim}(q_{embed},\{\mathbf{v}_{k}\mid m_{k}\in\mathcal{M}\}),k)$ (12)
Second, it defines expanded temporal windows $\mathbf{W}$ around the timestamps $(t_{s,k},t_{e,k})$ of these top-$k$ seed segments, using a temporal buffer $\delta$ to capture adjacent context:
$\mathbf{W}$ =$=\{[t_{s,k}-\delta,t_{e,k}+\delta]\mid k\in\mathbf{S}_{\text{query}}\}$ (13)
Third, the system retrieves information from the target modality (complementary to the query cue) by accessing segments $m_{j}$ that temporally overlap with any window in $\mathbf{W}$. This leverages temporal proximity for cross-modal association, retrieving, for example, visual details ($\mathbf{T}_{j}^{v},\mathbf{C}_{j}^{v}$) associated with an auditory cue within the same time frame:
$\mathbf{S}_{\text{target}}$ =$=\{\text{TargetInfo}(m_{j},\text{modality}_{\text{target}})\mid(t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}\}$ (14)
where $\text{TargetInfo}(m_{j},\text{modality}_{\text{target}})$ denotes extracting the specified target modality’s content (e.g., text, embeddings, pointers) from segment $m_{j}$, and the condition $(t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}$ is shorthand for overlapping with any interval $[t_{1},t_{2}]\in\mathbf{W}$. The hyperparameters $k$ and $\delta$ control the scope and temporal tolerance of this cross-modal retrieval mechanism.
4.3.3 Adaptive Reasoning for Final Answer Synthesis
Regardless of the retrieval path (Fast Retrieval or Detailed Recall), the collected evidence $r_{\text{retrieved}}$ must be synthesized into a final answer $a$. This is performed by the Adaptive Reasoning module $\Gamma$ (Figure 2b, “Adaptive Reasoning”), corresponding to the “AR” component (Table 1):
$a=\mathcal{R}_{\text{final}}(q)=\Gamma(q,r_{\text{retrieved}},\mathcal{E}_{\text{context}})$ (15)
$\Gamma$ is implemented using an LLM that receives query $q$, retrieved evidence $r_{\text{retrieved}}$, and context $\mathcal{E}_{\text{context}}$ with essential visual and auditory information from $r_{\text{retrieved}}$. The LLM synthesizes responses by analyzing evidence against the query, prioritizing relevance, resolving inconsistencies, and acknowledging limitations. This enables comprehensive responses that go beyond simple aggregation of the retrieved content.

### v2

@@ 4.1 Memory Formation
Our memory formation pipeline implements three computational primitives from the hippocampus: pattern separation (episodic segmentation), perceptual encoding, and memory consolidation (semantic abstraction). These stages transform continuous multimodal streams into a hierarchical memory structure with both detailed episodic traces and efficient semantic indices.
Episodic Segmentation (Pattern Separation / Threshold-based Boundary Detection). Following the DG’s pattern separation function that transforms overlapping inputs into distinct neural codes, we implement content-adaptive temporal segmentation to prevent interference between similar memories. Fixed-window chunking arbitrarily divides videos at regular intervals, often splitting coherent events mid-action or grouping unrelated scenes together. To preserve semantic integrity, our architecture segments the stream by detecting perceptual discontinuities that align with natural event boundaries. Specifically, we trigger a boundary at time $t$ when either visual similarity drops significantly between consecutive frames or level drops below a threshold indicating silence or scene breaks:
$\mathcal{S}_{t}=\text{1}[d_{v}(F_{t},F_{t-1})>\tau_{v}]\vee\text{1}[d_{a}(a_{t})>\tau_{a}],$ (1)
where visual dissimilarity $d_{v}$ is measured using Structural Similarity Index Measure (SSIM) (Eq. 2) with threshold $\tau_{v}$ implementing the sparsity constraint observed in DG neurons, and audio boundary detection $d_{a}$ (Eq. 3) uses threshold $\tau_{a}$ chosen through empirical validation to identify pauses in speech or sound that often coincide with activity transitions.
$d_{v}(F_{t},F_{t-1})$ $=1-\text{SSIM}(F_{t},F_{t-1}),$ (2)
$d_{a}(a_{t})$ $=-20\log_{10}\left(\sqrt{\frac{1}{N}\sum_{i=1}^{N}a_{i}^{2}}\right).$ (3)
These perceptually-motivated boundaries produce segments constrained between $t_{min}=5$ and $t_{max}=10$ seconds, aligning with human event segmentation timescales reported in prior work [40]. This content-aware segmentation produces a series of coherent temporal windows, $t_{s,i},t_{e,i}]$, forming the foundational units for memory encoding.
Perceptual Encoding. Mirroring how the entorhinal cortex provides multimodal input transformed through DG into CA3’s distributed representations, we employ specialized encoders that preserve both modality-specific features and cross-modal associations. Within each event segment, the system captures a rich, multimodal representation of the experience. This is not a monolithic encoding but a tripartite strategy that orchestrates specialized foundation models: a vision-language model generates dense textual descriptions ($\mathcal{T}_{v}$), a speech recognition model provides precise transcriptions ($\mathcal{T}_{a}$), and a cross-modal embedding model ($\mathcal{E}_{c}$) maps both visual frames and audio snippets into a shared feature space, enabling direct similarity comparisons across modalities. This distributed encoding mirrors CA3’s recurrent architecture where diverse inputs converge into unified episodic representations. These outputs are aggregated into a detailed episodic record, the ShortTermMemory object $m_{i}$:
$m_{i}=\{\mathbf{E}_{i},\mathbf{T}_{i},\mathbf{C}_{i},t_{s,i},t_{e,i}\},$ (4)
where $\mathbf{E}_{i}$ denotes the sequence of cross-modal embeddings from $\mathcal{E}_{c}$, $\mathbf{T}_{i}$ aggregates the textual outputs from both $\mathcal{T}_{v}$ (visual descriptions) and $\mathcal{T}_{a}$ (audio transcriptions), $\mathbf{C}_{i}$ contains pointers to the raw audiovisual data, and $t_{s,i},t_{e,i}$] marks the temporal boundaries of segment $i$.
Memory Consolidation (Synaptic Consolidation / Cosine-Similarity Sparse Filtering). Implementing synaptic and systems consolidation (with CA3/CA1 roles), our pipeline progressively abstracts from perceptual details to semantic summaries. Storing every fine-grained detail from a long stream is inefficient and leads to catastrophic interference between similar events. Our architecture addresses this with a consolidation mechanism that actively filters for novelty. This process filters redundant ShortTermMemory objects, retaining only those that introduce significant new information. For each segment $m_{i}$, a representative embedding $\mathbf{v}_{i}$ is computed by averaging the sequence of cross-modal embeddings $\mathbf{E}_{i}$ produced by the embedding model $\mathcal{E}_{c}$. The segment is retained in the consolidated set $K$ only if its embedding is sufficiently dissimilar from all previously stored memories, governed by a similarity threshold $\gamma$:
$K=\{i\mid\forall j\in K,j<i\implies\cos(\mathbf{v}_{i},\mathbf{v}_{j})<\gamma\}.$ (5)
This threshold $\gamma$ is chosen to encourage sparse retention, analogous to CA3 sparsity where only 2-5% of neurons are active for any given memory. This selective filtering creates an efficient memory store that is robust to distraction and optimized for retrieval. The consolidated set $K$ contains indices of retained segments. For clarity, we use $k\in K$ to denote consolidated segment indices and $m_{k}$ for their corresponding memory objects.
The cornerstone of our architecture’s efficiency is the creation of a high-level semantic index over the consolidated episodic records. This stage transforms detailed, fine-grained memories into abstract, language-based summaries, making vast stretches of time searchable at near-zero cost. For each consolidated memory $m_{k}$ where $k\in K$ denotes indices from the filtered set, an LLM ($\phi_{\text{LLM}}$) synthesizes its multimodal contents into a concise textual “gist,” $\mathbf{S}_{\theta_{k}}$:
$\mathbf{S}_{\theta_{k}}=\phi_{\text{LLM}}(\text{FormattedContext}(m_{k})),$ (6)
where $\text{FormattedContext}(\cdot)$ structures the multimodal content from $m_{k}$ (visual descriptions, transcriptions, and temporal information) into a prompt format suitable for the LLM. This summary becomes the centerpiece of a ThetaEvent object, our long-term memory data structure. The name “ThetaEvent” reflects hippocampal theta oscillations that bind experiences into coherent episodes. The ThetaEvent bundles the abstract summary with the representative embedding $\mathbf{v}_{k}$ and a pointer back to its detailed source memory, creating a powerful dual-representation system that bridges semantic abstraction and perceptual detail, precisely the function performed by CA1 in biological memory consolidation.
$\theta_{k}=\{\mathbf{v}_{k},\mathbf{S}_{\theta_{k}},\mathbf{T}_{\theta_{k}},\mathbf{I}_{\theta_{k}},\mathbf{A}_{\theta_{k}}\},$ (7)
where $\mathbf{v}_{k}$ is the representative embedding, $\mathbf{S}_{\theta_{k}}$ is the generated semantic summary, $\mathbf{T}_{\theta_{k}}$ contains temporal references linking to the source $m_{k}$, and $\mathbf{I}_{\theta_{k}}$, $\mathbf{A}_{\theta_{k}}$ store key visual and auditory context used in summary generation.
4.2 Hierarchical Memory Retrieval (Pattern Completion / Confidence-Gated Two-Path Lookup)
The hippocampus supports two complementary retrieval modes: rapid semantic access through CA1 and detailed pattern completion through CA3 recurrence. Our retrieval architecture directly implements this biological duality, resolving the critical trade-off between speed and fidelity by dynamically choosing its search strategy based on the query’s demands, facilitating complex forms of reasoning such as associative recall and cross-modal pattern completion.
Query-Driven Pathway Selection. Implementing the hippocampus’s dual retrieval modes: CA1’s rapid semantic access and CA3’s detailed pattern completion, we employ confidence-gated hierarchical retrieval. Upon receiving a query $q$, the system first attempts Fast Retrieval ($\Phi_{\text{fast}}$), implementing CA1’s semantic pathway by searching exclusively on the abstract ThetaEvent summaries. This path rapidly resolves high-level and gist-based queries. Only if the confidence of this initial result falls below a tuned threshold $\tau$ does the system escalate to the more resource-intensive Detailed Recall ($\Psi_{\text{detailed}}$), which engages CA3’s pattern completion mechanisms. This confidence score is generated by the LLM’s internal calibration when attempting to answer from summaries alone, reflecting its assessment of whether sufficient information exists in the abstract representations. This on-demand escalation provides a “zoom-in” capability, accessing the rich, consolidated ShortTermMemory objects for evidence-grounded reasoning.
$\mathcal{R}(q)=\begin{cases}\Phi_{\text{fast}}(\Theta,q)&\text{if }\text{conf}(\Phi_{\text{fast}})>\tau,\\ \Psi_{\text{detailed}}(q,\mathcal{M},\mathcal{Q}_{\text{type}}(q))&\text{otherwise},\end{cases}$ (8)
where $\Theta=\{\theta_{k}\mid k\in K\}$ denotes the collection of ThetaEvent objects.
Detailed Recall with Cross-Modal Pattern Completion. CA3’s recurrent connectivity enables pattern completion: reconstructing complete memories from partial cues. This biological computation is most powerful for cross-modal binding, where a cue in one modality triggers recall of associated information in another modality. When detailed recall is invoked, the system deploys a specialized search strategy tailored to the query’s modality. For single-modality queries, it performs either an embedding-based Feature Search or a text-based Semantic Search. The key innovation, however, is its mechanism for cross-modal associative recall, which uses temporal co-occurrence as a powerful proxy for episodic binding, mirroring how CA3 neurons bind coincident inputs through synaptic plasticity.
This process begins by using the query cue to find relevant seed segments in the target memory store:
$\mathbf{S}_{\text{query}}=\text{TopK}(\text{sim}(q_{embed},\{\mathbf{v}_{k}\mid m_{k}\in\mathcal{M}\}),k).$ (9)
It then defines expanded temporal windows $\mathbf{W}$ around these seeds to capture the surrounding event context, implementing the temporal receptive fields of hippocampal place cells.
$\mathbf{W}=\{[t_{s,k}-\delta,t_{e,k}+\delta]\mid k\in\mathbf{S}_{\text{query}}\}.$ (10)
Finally, it retrieves information from the target modality by searching for any memory segments that temporally overlap with these windows, effectively reconstructing the integrated experience.
$\mathbf{S}_{\text{target}}=\{\text{Extract}_{\text{target}}(m_{j})\mid\\ (t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}\},$ (11)
where $\text{Extract}_{\text{target}}(m_{j})$ extracts modality-specific content from segment $m_{j}$: visual descriptions from $\mathbf{T}_{j}$ for visual targets, or transcriptions for auditory targets. We denote segment $j$’s temporal boundaries as $(t_{s,j},t_{e,j})$.
To illustrate cross-modal pattern completion, consider the query: “What was on the screen when the applause started?” The system identifies the auditory cue (“applause”) and visual target (“on the screen”), with low confidence from Fast Retrieval triggering Detailed Recall due to the specific temporal link. Following Eq. 9, it finds seed segment $m_{k}$ containing the applause sound, then forms temporal window $\mathbf{W}$ around it (Eq. 10)—implementing how hippocampal neurons encode temporally extended episodes. Visual information from segments overlapping $\mathbf{W}$ is retrieved (Eq. 11) and passed to Adaptive Reasoning for synthesis.
@@ 4.2 Hierarchical Memory Retrieval (Pattern Completion / Confidence-Gated Two-Path Lookup)
Table 1: Performance comparison of HippoMM against existing methods and ablation variants on HippoVlog. PT: Processing Time; ART: Average Response Time; A+V: Cross-modal (Audio+Visual) accuracy; A: Audio-only accuracy; V: Visual-only accuracy; S: Semantic understanding accuracy. Best results are in bold, second best are underlined. HippoMM significantly outperforms prior methods in all modality-specific tasks while maintaining efficient processing and response times. Ablation studies demonstrate the importance of each component: Detailed Recall (DR), Fast Retrieval (FR), and Adaptive Reasoning (AR).
@@
NotebookLM [11] 28.40% 23.20% 28.00% 26.80% 26.60%
VideoRAG [28] 9.46h 112.5s 63.6% 67.2% 41.2% 84.8% 64.2%
Ola [21] 79.4s 72.4% 85.6% 57.6% 84.0% 74.9%
VideoLLaMA 3 [41] 58.3s - - 70.8% 75.2% 73.0%
Simple-Assoc-Memory (SAM) 32.0% 32.8% 33.2% 23.2% 30.3%
Gemini-3-Flash [10] 44.0% 45.2% 40.0% 56.0% 46.3%
GPT-5 [25] 72.0% 73.2% 45.6% 88.0% 69.7%
Qwen2.5-Omni [37] 71.2% 69.6% 51.2% 73.6% 66.4%
@@
HippoMM w/o DR 5.09h 6.39s 39.2% 65.6% 48.0% 92.0% 61.2%
HippoMM w/o FR 5.09h 19.54s 72.0% 79.6% 63.6% 83.2% 74.6%
HippoMM (Qwen2.5-14B) 5.09h 15.7s 69.3% 77.1% 56.6% 76.4% 70.8%
HippoMM (EOR-only) 5.09h 63.2% 70.4% 60.8% 90.0% 71.1%
HippoMM ($\gamma=0.80$) 5.09h 20.4s 69.2% 78.4% 64.0% 90.8% 75.6%
HippoMM ($\gamma=0.90$) 5.09h 20.4s 71.6% 78.8% 64.4% 90.0% 76.2%
@@
Adaptive Reasoning. While the hippocampus provides memory access, the prefrontal cortex performs executive control and conflict resolution. Our Adaptive Reasoning module implements this cognitive function, particularly crucial when reconciling outputs from the dual retrieval pathways. The module $\Gamma$ serves as a locus for cognitive reflection, especially when the system escalates from Fast Retrieval to Detailed Recall. In such cases, $\Gamma$ receives both summary-based evidence from $\Phi_{\text{fast}}$ and fine-grained evidence from $\Psi_{\text{detailed}}$, reconciling potentially conflicting accounts. For instance, if a summary suggests general activity while detailed recall reveals specific contradictory actions, the module prioritizes high-fidelity, grounded evidence from the detailed pathway—mirroring prefrontal-hippocampal interactions where top-down control resolves memory conflicts. The module synthesizes a final answer that accurately represents the most reliable evidence:
$a=\mathcal{R}_{\text{final}}(q)=\Gamma(q,r_{\text{retrieved}},\mathcal{E}_{\text{context}}).$ (12)
By integrating this reflective step, the architecture ensures its output is not just retrieved facts but a reasoned interpretation, demonstrating that cognitive architectures provide both biological plausibility and functional superiority.